### v1

# #1#1#1# 1 (1)

Abstract
Long-context reasoning requires accurately identifying relevant information in extensive, noisy input contexts. Previous research shows that using test-time learning to encode context directly into model parameters can effectively enable reasoning over noisy information. However, meta-learning methods for enabling test-time learning are prohibitively memory-intensive, preventing their application to long context settings. In this work, we propose PERK (Parameter Efficient Reasoning over Knowledge), a scalable approach for learning to encode long input contexts using gradient updates to a lightweight model adapter at test time. Specifically, PERK employs two nested optimization loops in a meta-training phase. The inner loop rapidly encodes contexts into a low-rank adapter (LoRA) that serves as a parameter-efficient memory module for the base model. Concurrently, the outer loop learns to use the updated adapter to accurately recall and reason over relevant information from the encoded long context. Our evaluations on several long-context reasoning tasks show that PERK significantly outperforms the standard prompt-based long-context baseline, achieving average absolute performance gains of up to 90% for smaller models (GPT-2) and up to 27% for our largest evaluated model, Qwen-2.5-0.5B. In general, PERK is more robust to reasoning complexity, length extrapolation, and the locations of relevant information in contexts. Finally, we show that while PERK is memory-intensive during training, it scales more efficiently at inference time than prompt-based long-context inference.
1 Introduction
Large language models (LLMs) struggle with identifying and reasoning over relevant information in long, potentially noisy, input contexts (i.e., long-context reasoning [2, 34, 36, 38, 65]), including modern models with substantial context windows [32, 35, 23] such as GPT-4 (claimed context window of 128k tokens [23]). These challenges are not surprising, as longer contexts contain more irrelevant distractor information, and may also reflect more complex reasoning problems (i.e., multiple hops), both of which increase the challenge of reasoning tasks and degrade LLM performance [13]. Furthermore, positional biases in long-context models focus disproportionate attention on the beginning or end of contexts, neglecting intermediate information [39].
In this work, we address these challenges of long-context reasoning by reframing it as test-time learning, enabling models to encode context at inference time using gradient-based parameter adaptation. Our approach processes and compresses lengthy sequences as a batch (or multiple batches) of shorter segments of the long context, rather than processing the full sequence token-by-token. The updated model can then use the newly stored parametric information (and its initial parametric knowledge) to answer relevant questions. During training, we operate two nested optimization loops: the inner loop, which is also used at test time, encodes the long context segment batches into model parameters, while the outer loop learns to retrieve and reason over relevant information from the updated parameters that now encode the long context.
However, meta-learning algorithms for training test-time learning methods entail prohibitive memory overhead [17, 13, 25], particularly when updating the complete set of model parameters, because they require backpropagating through a complete multi-step optimization trajectory. To reduce the memory overhead, we introduce Parameter-Efficient Reasoning over Knowledge (PERK), a lighter-weight meta-learning algorithm that keeps training-time memory costs low by: (1) internalizing the context into a Low-Rank Adapter (LoRA [24]), and (2) backpropagating the outer loop over only a truncated trajectory from the final few inner-loop adaptation steps [54]. By lowering the memory cost, PERK scales more efficiently at training, supporting larger models and longer contexts.
We benchmark PERK relative to models finetuned for classical in-context reasoning (FT-ICR) on multiple Needle-in-a-Haystack benchmarks, demonstrating PERK’s advantages. A small GPT-2 model (with 127M parameters and a native 1k context window) can be scaled by PERK to reason effectively over 8k token contexts, achieving a 60% absolute performance gain over fine-tuning GPT-2 for in-context reasoning. For larger models (Qwen-2.5-0.5B) with a native 32k context window, PERK surpasses FT-ICR by a 27% absolute improvement. As relevant facts in Needle-in-a-Haystack problems are often easily distinguishable from distractor text, we also propose a more challenging evaluation setting where relevant facts are surrounded by distributionally-similar distractor facts (i.e., Drops-in-the-Ocean). In this more complex setting, PERK models again outperform all baselines, including FT-ICR across the GPT-2, Qwen, and Mamba model families.
Further analyses reveal PERK’s strong test-time length extrapolation ability. Models trained with PERK on 1k-token contexts successfully extrapolate performance to unseen 32k-token contexts with only a 42% degradation, far better than FT-ICR’s 52% decline. In addition, unlike FT-ICR’s dramatic drops in performance (up to 90%) when relevant information shifts positions at test time, PERK robustly generalizes across contexts where relevant information is variably distributed along context positions. Finally, our inference complexity analysis shows that PERK’s memory footprint scales more efficiently than FT-ICR, as does its wall-clock runtime on extremely long contexts (> 32k tokens).
2 Background and Preliminaries
Causal Language Models   In this work, we focus on causal language models (CLM) $f_{\bm{\theta}}$ with a decoder-only transformer [51] architecture $f$ and parameters ${\bm{\theta}}$. The common objective in autoregressive-style training of a CLM is the token-level negative log-likelihood (NLL). For a collection of text sequences ${\mathcal{D}}=\{x\}$ with total length $N_{\mathcal{D}}$, this is defined as:
$\mathcal{L}_{\text{NLL}}({\bm{\theta}},{\mathcal{D}})\;=\;-\;\frac{1}{N_{\mathcal{D}}}\sum_{x\in{\mathcal{D}}}\;\;\sum_{t=1}^{\mathcal{T}}\log p_{\bm{\theta}}\bigl{(}x_{t}\;\big{|}\;{\bm{x}}_{<t}\bigr{)}.$ (2)
For training, ${\bm{\theta}}$ is optimized iteratively via an optimization algorithm $\mathcal{A}lg:(\mathcal{L},{\mathcal{D}},{\bm{\theta}},\bm{h})\mapsto{\bm{\theta}}^{\prime}$ where $\mathcal{L}$ is the objective function, ${\mathcal{D}}$ is the training corpus, ${\bm{\theta}}$ are the initial parameters, ${\bm{\theta}}^{\prime}$ are the obtained parameter values, and $\bm{h}$ are the hyperparameters, such as optimizer choice and training length.
Reasoning with CLMs   We denote reasoning problems as tuples $r=({\mathcal{K}},{\bm{q}},y)$ of context ${\mathcal{K}}$, question ${\bm{q}}$, and target response $y$. With the pre-trained large CLMs’s ability to reason over facts in input contexts with reasonable success [14], prompt-based in-context reasoning (ICR) has become popular for automating reasoning tasks. In this setting, ${\mathcal{K}}$ and ${\bm{q}}$ are combined via a template into a single prompt ${\bm{x}}(r)$. The model’s parameters are optimized to predict the tokens of the correct answer:
$\mathcal{L}_{\text{reason}}({\bm{\theta}},{\mathcal{D}})\;=\;-\;\frac{1}{N_{{\mathcal{D}}_{y}}}\sum_{r=({\mathcal{K}},{\bm{q}},y)\in{\mathcal{D}}}\;\sum_{t=1}^{\mathcal{T}_{y}}\log p_{\bm{\theta}}\bigl{(}y_{t}\,\big{|}\,{\bm{x}}(r),y_{<t}\bigr{)}$ (3)
where ${\mathcal{D}}$ is the training corpus, $\mathcal{T}_{y}$ is the length of an example answer, and $N_{{\mathcal{D}}_{y}}$ is the total length of all answers in the dataset.
Reasoning as Test-time Learning   In this setting, models simulate reasoning by first encoding the given context ${\mathcal{K}}$ into a CLM’s parameters ${\bm{\theta}}$, and then using the updated parameters $\phi({\bm{\theta}},{\mathcal{K}})$ to generate a response to the given question ${\bm{q}}$ [13]. The context is encoded by adapting the model’s parameters using a CLM objective over the context: $\phi({\bm{\theta}},{\mathcal{K}})=\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},{\bm{\theta}},\bm{h})$. We then use $f_{\phi}$ to generate answers for any relevant questions ${\bm{q}}$.
Test-time learning methods for reasoning stem from a key observation: CLMs can store and retrieve diverse information in their parameters from pre-training [52, 5]. Yet these same models struggle with long-context problems [23, 32, 65, 35], despite the contexts containing far less information than what they’ve seen in pre-training. CLMs may reason better with knowledge in their parameters than in context, motivating methods that encode new context parametrically for reasoning.
Training a CLM for Test-time Learning   Given a distribution $R$ of possible reasoning problems, the optimal parameters ${\bm{\theta}}^{*}$ for a test-time learning CLM minimize the expected reasoning loss of the CLM’s adaptation to each reasoning problem $r\sim R$:
${\bm{\theta}}^{*}=\operatorname*{arg\,min}_{{\bm{\theta}}}\;\mathop{\mathbb{E}}_{r\sim R}\bigg{[}\mathcal{L}_{\text{reason}}\Bigl{(}\phi({\bm{\theta}},{\mathcal{K}}),\{({\bm{q}},y)\}\Bigr{)}\bigg{]}$
where $\phi({\bm{\theta}},{\mathcal{K}})$ is a CLM adaptation of ${\bm{\theta}}$ to the context ${\mathcal{K}}$: $\phi({\bm{\theta}},{\mathcal{K}})=\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},{\bm{\theta}},\bm{h})$.
Meta-learning algorithms, such as MAML [17], can be used for training effective test-time learning methods. In MAML, we obtain a meta-model $f_{{\bm{\theta}}}$ via bi-level optimization: the outer loop samples batches of reasoning problems from a training dataset of reasoning problems, the inner loop obtains the adaptations of ${\bm{\theta}}$ to each problem’s context ${\mathcal{K}}$, and the outer loop then optimizes the meta-parameters based on the performance of each $f_{\phi({\bm{\theta}},{\mathcal{K}})}$ on the corresponding questions. The outer loop optimization backpropagates through the entire inner loop, an operation that becomes increasingly expensive with more parameters and inner loop update steps. This approach—adapting a model with a CLM objective on context in the inner loop while meta-optimizing a reasoning objective in the outer loop—was first introduced by Reckoning [13], which demonstrated its potential.
3 PERK: Meta-Learning Parameter-Efficient Reasoning over Knowledge
While prior work has demonstrated the potential of test-time learning as a reasoning paradigm [13], these approaches are hindered by high memory overhead during training. The underlying meta-learning approach, MAML [17], requires backpropagating through the expected test-time learning phase (the inner loop). This gradient unrolling process involves computing higher-order derivatives, which are known to have substantial memory requirements [57], thereby limiting the number of model parameters that can be optimized and the effective sequence length that can be encoded.
To scale the training of larger and more capable test-time learning models, we propose PERK  (Parameter Efficient Reasoning over Knowledge), which uses a parameter-efficient test-time adaptation to reduce both the size and amount of gradient unrolling required during optimization. Specifically, in the inner and outer loops, we only optimize a parameter-efficient adapter, LoRA [24], to encode the contexts, and we apply truncated gradient unrolling [54] to reduce the memory cost of backpropagating through the inner loop. We visualize the PERK training process in Figure 1.
Figure 1: Meta-learning PERK for long-context reasoning. The training procedure involves a nested inner and outer loop. The inner loop optimizes the likelihood of a batch of long context segments with respect to the parameters of the LoRA-based memory scratchpad. In the outer loop, the model uses the encoded information in the memory scratchpad to answer questions. In both cases, only the memory scratchpad parameters are updated while the base LLM’s parameters are frozen.
3.1 Parameter-Efficient Meta-Learning with Low-Rank Adaptation
Similar to Reckoning [13], PERK uses bi-level optimization to prepare a model for reasoning via test-time adaptation to any given context. At inference time, a model is adapted to the context via CLM optimization on that context, after which the adapted model is used to answer any relevant questions. However, unlike Reckoning, which directly finetunes the full model at training and test time, PERK only adapts a low-rank adapter [24]. During training, PERK encodes a context into the adapter through gradient-based updates, leaving the base model parameters unchanged. We denote by ${\bm{\theta}}_{\text{base}}$ and ${\bm{\theta}}_{\text{adapter}}$ the parameters of the base model and the LoRA adapter. ${\bm{\theta}}_{\text{base}}$ is never optimized, while ${\bm{\theta}}_{\text{adapter}}$ is meta-learned to a good initial state for its eventual test-time adaptations.
Test-Time Learning and Inner Loop Training    We set the test time learning objective (thus also the inner loop objective) to causal language modeling of the context: $\mathcal{L}_{\text{NLL}}({\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}))$, and set the adaptation algorithm $\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),\bm{h})$ to only update ${\bm{\theta}}_{\text{adapter}}$. The adaptation algorithm computes the gradient $\nabla\mathcal{L}_{\text{NLL}}$ on a batch of sub-sequences from the full context sequence ${\mathcal{K}}$. This compression as a parallel batch (or multiple batches) allows us to process lengthy sequences beyond the model’s context window. The exact manner by which ${\mathcal{K}}$ is broken into batches is a hyperparameter of the algorithm, and for efficiency reasons, we set it to one batch during training.
Outer Loop Training    In the outer loop, we optimize the meta parameters ${\bm{\theta}}_{\text{adapter}}$ to learn from $R$, the corresponding distribution of reasoning problems. The optimal ${\bm{\theta}}_{\text{adapter}}$ minimize the expected reasoning loss of the adapter’s adaptation to each reasoning problem $r=({\mathcal{K}},{\bm{q}},y)\sim R$:
${\bm{\theta}}_{\text{adapter}}^{*}=\operatorname*{arg\,min}_{{\bm{\theta}}_{\text{adapter}}}\;\mathop{\mathbb{E}}_{r\sim R}\bigg{[}\mathcal{L}_{\text{reason}}\Bigl{(}{\bm{\theta}}_{\text{base}},\,\phi_{\text{adapter}}(({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),{\mathcal{K}}),\,\{({\bm{q}},y)\}\Bigr{)}\bigg{]},$ (4)
where $\phi_{\text{adapter}}(({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),{\mathcal{K}})=\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),\bm{h})$ is the parameter-efficient CLM adaptation of ${\bm{\theta}}_{\text{adapter}}$ to the context ${\mathcal{K}}$ obtained via the test-time adaptation. We may use any optimization algorithm to optimize ${\bm{\theta}}_{\text{adapter}}$, and in our experiments, we employ AdamW [40].
3.2 Scalable Meta-Learning with Truncated Gradient Unrolling
To optimize Equation (4) with gradient-based methods, we need to differentiate through $\mathcal{A}lg$, which involves higher-order derivatives and saving the complete trajectory to compute the meta-gradient (i.e., the gradient of $\mathcal{L}_{\text{reason}}$). However, this creates high memory costs, capping the method’s applicability. To reduce memory costs, we truncate the backpropagation for the meta-gradient computation. We briefly describe this approach here, presenting the case where $\mathcal{A}lg$ optimizes ${\bm{\theta}}_{\text{adapter}}$ via direct gradient descent for simplicity (though we use more complicated optimizers for the inner loop in practice).
Gradient Unrolling (GU)    We denote an N-step inner-loop optimization’s intermediate states by:
$\phi_{\text{adapter}}^{(0)}\;=\;{\bm{\theta}}_{\text{adapter}},\quad\phi_{\text{adapter}}^{(n+1)}=\phi_{\text{adapter}}^{(n)}-\alpha\,g^{(n)},\quad\phi_{\text{adapter}}^{*}\;=\;\phi_{\text{adapter}}^{(N)},$
where $g^{(n)}\;=\;\nabla_{\phi_{\text{adapter}}^{(n)}}\!\mathcal{L}_{\text{NLL}}\bigl{(}{\bm{\theta}}_{\text{base}},\,\phi_{\text{adapter}}^{(n)},\,{\mathcal{K}}\bigr{)}$ is the gradient at step $n$, and $\alpha$ is the learning rate.
In vanilla MAML, to compute the gradient of the outer-loop loss $\mathcal{L}_{\text{reason}}$ with respect to the meta-parameters ${\bm{\theta}}_{\text{adapter}}$, one backpropagates through all $N$ inner steps. By the chain rule, this equates to accumulating the Jacobian $J^{(n)}=\frac{\partial\phi_{\text{adapter}}^{(n+1)}}{\partial\phi_{\text{adapter}}^{(n)}}$ from each step:
$\nabla_{{\bm{\theta}}_{\text{adapter}}}\mathcal{L}_{\text{reason}}\;=\;\frac{\partial\mathcal{L}_{\text{reason}}}{\partial\phi_{\text{adapter}}^{(N)}}\;J^{(N-1)}J^{(N-2)}\dotsm J^{(0)}\;\frac{\partial\phi_{\text{adapter}}^{(0)}}{\partial{\bm{\theta}}_{\text{adapter}}}\;=\;\frac{\partial\mathcal{L}_{\text{reason}}}{\partial\phi_{\text{adapter}}^{(N)}}\;\prod_{n=0}^{N-1}\!J^{(n)},$ (5)
where $J^{(n)}$ can be computed via: $J^{(n)}=I-\alpha\,H^{(n)},\;H^{(n)}=\nabla^{2}_{\phi_{\text{adapter}}^{(n)}}\!\mathcal{L}_{\text{NLL}}.$ Each Jacobian $J^{(n)}$ requires storing the computation graph of step $n$ in memory, so the cost of the backward pass scales linearly with $N$, creating a high memory requirement for gradient unrolling.
Truncated Gradient Unrolling (TGU)    To reduce memory, we follow [54] and run the inner loop for all $N$ specified update steps, but only store the computational graph for the last $T\ll N$ steps. This substantially reduces the computational memory requirements of the method, at the cost of slightly less accurate meta-gradients for the outer-loop optimization of ${\bm{\theta}}_{\text{adapter}}$. To implement this approximation, we do not retain the inner loop optimization’s computation graph until the start of the retention window: the last $T$ inner-loop optimization steps. Since all Jacobians before step $n_{\text{start}}$ are treated as constants, the Jacobian product in Equation (5) is effectively cut:
[…]
4 Experiments
In this section, we empirically evaluate PERK’s long-context reasoning capabilities in two challenging scenarios: reasoning over (1) Needles-in-a-Haystack (NIAH, [32]) and (2) Drops-in-an-Ocean (DIO), a novel long context evaluation setting we propose.
Baselines   We compare our method to two long context reasoning adaptations applied to multiple models. In particular, we finetune transformer [60] and Mamba [21] models for prompting-based in-context reasoning (FT-ICR), as in conventional long-context finetuning [18, 19]. We apply FT-ICR to GPT-2 (127M parameters) [51], a small model with a limited context window (1024 tokens) and no long-context pretraining, Qwen-2.5-0.5B [50] (with a 32k token context window), and to two Mamba [21] models with sizes of 130M and 1.4B parameters. We refer to these baselines as FT-ICR (GPT-2), FT-ICR (Qwen), FT-ICR (Mamba-130M), and FT-ICR (Mamba-1.4B). Whenever evaluating tasks with contexts longer than the 1024 token limit of GPT-2, in order to effectively fine-tune it, we first provide GPT-2 with a series of additional positional embeddings instantiated using the interpolation method proposed in [29]. The Qwen models, meanwhile, all have sufficient context length for our experiments, and the Mamba models are designed to work with sequences of any length. In our experiments, we refer to models trained with PERK as PERK (GPT-2) and PERK (Qwen) depending on their base model. With only the LoRA parameters being trainable, PERK (GPT-2) contains 38M trainable parameters, and PERK (Qwen) has 140M trainable parameters.
4.1 Reasoning Over Needles-in-a-Haystack
Figure 2: Performance on NIAH with BabiLong and DIO with Student Records. All models are trained and tested on sequences ranging from 1K to 8K tokens. All PERK models are trained to first generate the supporting facts relevant to a query, followed by the final answer prediction. In contrast, baseline models directly generate the answer, as this approach yields better performance for them in this setting. The number of trainable parameters for each method is indicated in the legend.
Setup
We first evaluate PERK on reasoning over Needles-in-a-Haystack using the BabiLong framework [32] for assessing models’ reasoning abilities across relevant and distractor facts scattered in very long documents. We mainly focus on tasks involving single-hop (QA1), two-hop (QA2), and three-hop (QA3) reasoning. We use the framework to construct the tasks with long natural documents sampled from FineWeb-Edu [41], ensuring the content postdates our base model’s knowledge cutoff.
Results
In the top row of Figure 2, we show the evaluation results on BabiLong. PERK outperforms all baselines in all cases we test (across all models), including when compared to larger models. In particular, PERK (GPT-2) outperforms even FT-ICR (Qwen) and FT-ICR (Mamba-1.4B) for all task complexities and context lengths, with an average 20% performance gain. PERK (Qwen) achieves the highest accuracy across all task complexities (QA1, QA2, QA3) and context lengths (1K to 8K). PERK is also more robust to increasing context length and task complexity than the baselines. At best, PERK (Qwen) loses almost no accuracy with growing context length (for QA1 and QA2), and at worst (for QA3), it loses ${\sim}10\%$. When averaging across context lengths, PERK (Qwen)’s performance gain over FT-ICR (Qwen) increases from 10% to 41% when the number of reasoning hops increases from 1-hop (QA1) to 3-hop (QA3).
4.2 Reasoning Over Drops-in-the-Ocean
Needle-in-a-Haystack (NIAH) datasets for long-context evaluation typically have a fundamental limitation: the target information often exhibits stylistic differences from surrounding irrelevant text, creating an artificially simple test for identifying relevant information. In practical scenarios, relevant and irrelevant information will likely be distributionally similar, increasing the difficulty of identifying critical facts. As a result, we propose Drops-in-the-Ocean (DIO), a new evaluation setting that forms long contexts from structurally similar documents.
Setup
We construct a synthetic dataset, Student Records, where each context simulates a database containing multiple student records. Each record includes several attributes: ID, name, school, major, and grade. Unique IDs and names are generated for each database instance. Schools and majors are sampled with replacement from predefined pools, and grades are integers sampled uniformly from the range [0, 100]. Context length scales directly with the number of student records included. Inspired by BabiLong, we define evaluation tasks of varying complexity: Recall (retrieving attributes for a specific ID), Relation (comparing attributes between two IDs), and Aggregate (calculating the maximum, minimum, and average grade across all students). Crucially, to test generalization, specific entities (e.g., IDs, names) and attribute value combinations encountered in the test set are disjoint from those used during training. Examples are available in the Appendix.
Results
Results on the Student Records in Figure 2 mirror the trends seen in BabiLong. FT-ICR (GPT-2) fails to reason in the DIO setting with 0% accuracy or random performance (for Relation, the random performance is 50%). FT-ICR (Qwen) appears to be the strongest baseline, especially on recalling with a 1K context, where it briefly matches PERK (Qwen)’s near-perfect accuracy. PERK models again outperform all baselines. PERK (Qwen) and PERK (GPT-2) consistently achieve high accuracy (often >85%) and are more robust to increasing context length, even on the more challenging aggregation task with the 8k length, which requires the model to aggregate information across all 8k tokens. With Qwen as a base model, PERK achieves a 9% average gain over FT-ICR.
5 Analysis
5.1 Test-Time Length Generalization
We evaluate PERK’s ability to generalize to longer inputs than those seen in training [48].
Setup
We evaluate using the BabiLong tasks, which allow us to generate contexts of arbitrarily long lengths, and train a Qwen-2.5-0.5B base model using PERK on fixed-length contexts ranging from 1K to 8K tokens (as well as an FT-ICR baseline). Subsequently, we evaluate each model on contexts of up to 32K tokens, measuring each approach’s ability to generalize to new, longer lengths (i.e., extrapolation), as well as shorter contexts (i.e., interpolation). Overall, we train on contexts of increasing lengths (1K to 8K tokens) and test on contexts ranging from 1K to 32K tokens.
PERK generalizes to new context lengths at test time.
Figure 5 shows that PERK outperforms FT-ICR on the BabiLong QA1 (single-hop) and QA2 (two-hop) tasks for all generalization settings. For extrapolation, FT-ICR’s performance drops drastically as the inference context length exceeds the training context length, especially for shorter training contexts.
Figure 3: Test-time length extrapolation beyond 32K on BabiLong QA1 (a) and QA2 (b). Both PERK and FT-ICR are trained on 8K-token sequences. The context length for inference grows from 64K to 128K. PERK extrapolates substantially better than FT-ICR.
While PERK’s accuracy also decreases (from a higher starting point), the degradation is dampened at longer test lengths. For QA1 (Figure 5a), PERK trained on 1K sequences shows only a 42% drop on 32K sequences while FT-ICR exhibits a 52% decline. When trained on 8K sequences, PERK drops only 5% on 32K sequences, while FT-ICR still exhibits a 32% drop. Similar dynamics are observed on the more challenging QA2 task (Figure 5b). To stress-test the limit of PERK models’ length extrapolation ability, we test on sequences with context lengths beyond the context window of the Qwen-2.5 model – 32K tokens in Figure 3. Specifically, we test on sequences with tokens increasing from 64K to 128K, and demonstrate that PERK consistently extrapolates to longer contexts than the FT-ICR baseline. We acknowledge that with 128K token context windows, PERK’s accuracy experiences a noticeable drop to 61.4% for QA1 accuracy and 44.4% for QA2 accuracy. However, these scores remain substantially better than the 0% QA1 and QA2 absolute performance of the FT-ICR baselines at this same context length.
Meanwhile, for interpolation, FT-ICR performance drops on shorter contexts when finetuned on longer ones, matching observations from prior work [19]. In contrast, PERK maintains its performance and even increases performance for the QA2 task. Overall, PERK demonstrates strong length generalization, extrapolating to longer contexts while preserving high performance on shorter ones.
Figure 4: Test-time context length generalization evaluation on BabiLong QA1 (a) and QA2 (b): comparison between PERK and FT-ICR on the Qwen-2.5-0.5B model. The y-axis represents the training context lengths, while the x-axis indicates various test-time context lengths. We test for both interpolation (test lengths shorter than the training length) and extrapolation (test lengths longer than the training length). Bordered cells denote the boundary: evaluation on context lengths equal to those in training. PERK shows stronger generalization across both settings.
Figure 5: Positional Bias. Comparison of PERK and FT-ICR on 4K and 8K contexts, on Qwen-2.5-0.5B. We train on problems where the relevant information appears in the beginning (Pre), middle (Mid), or end (Post) of the context, and evaluate on all three positional settings. We also train models on contexts where the relevant information is randomly located (Rnd), testing these on all four positional distributions (Pre, Post, Mid, Rnd). Bordered cells show in-distribution performances. PERK demonstrates strong positional robustness.
5.2 Robustness to Positional Biases
Prior work has shown that the position of relevant information in long contexts affects a language model’s ability to utilize that information [39]. In this experiment, we test PERK’s robustness to positional biases by evaluating its test-time performance on contexts where the position of relevant information is distributed differently from those seen in training.
Setup
We use API documents from [45], where each document has an associated API call, and create long contexts containing multiple API-call-to-document pairs. In this task, the model must retrieve the correct API call based on user instructions and documents. Using fixed 4k and 8k token contexts, we train PERK and FT-ICR models on contexts with the target document at varying positions and test them on contexts distributed across all positions.
PERK generalizes regardless of information position in the context.
Figure 5 compares PERK to FT-ICR under different position distributions (at train and test time) and context lengths. When trained on contexts with relevant documents randomly located (Rnd) throughout the context, PERK outperforms FT-ICR by 24%. Furthermore, FT-ICR shows large performance drops when the relevant document consistently appears at different locations (Pre, Mid, and Post) at test time. Meanwhile, shifting relevant positions at test-time shows minimal effect (within 1-2%) on PERK.
To further stress-test positional generalization, we also force relevant documents into specific positions during training and testing, spanning the full context. We find that FT-ICR easily overfits to the position pattern, completely failing to generalize to test-time position changes (performance drops to close to 0% when the position shifts at test time).
Although PERK underperforms FT-ICR when the train-test positions are the same and the beginning (Pre) or end (Post) of the context, performance remains consistent regardless of the positional distribution shift between train and test time. We attribute PERK’s robustness in this setting to the fact that documents in the long context are encoded as members of a batch into the parameter space, limiting the overall impact of absolute position in the sequence. Meanwhile, FT-ICR directly aligns specific positions with answers during training.
Figure 6: (a): Peak GPU memory usage during training with context lengths ranging from 1K to 8K tokens for Reckoning and PERK. While PERK successfully scales to 8K tokens, Reckoning encounters out-of-memory (OOM) errors at shorter context lengths. (b): Memory footprint and wall-clock runtime during inference as context length increases (up to 128K tokens), comparing PERK with FT-ICR. Curves that terminate before 128K indicate that the method failed with an OOM error at longer context lengths, preventing further measurement. PERK demonstrates more efficient scaling in both memory and runtime, particularly for extremely long sequences.
5.3 Training and Inference Complexity Analysis
We first show PERK’s training scalability compared to prior test-time learning method Reckoning [13] by measuring each method’s peak training memory across different context lengths. We then evaluate PERK’s training efficiency on long contexts by measuring its inference memory cost and run-time, compared to in-context reasoning with finetuned models.
Setup
We complete all runs on a single Nvidia H100 GPU with 93 GB of VRAM. We perform four gradient update steps for the test-time and inner loop adaptations. For training, we have four truncation levels: truncate 0 (basically gradient unrolling without truncation), 1, 2, and 3 steps. For inference, both methods generate 64 tokens using greedy decoding with Huggingface’s text generation pipeline. We measure runtime complexity in seconds (wall-clock time) and memory complexity in gigabytes. Both methods undergo 10 hardware warm-up iterations.
As explained in Section 3, PERK splits long sequences into batches of shorter sequences. Here, each batch has an effective context length defined as 128 tokens. Since PERK performs gradient-based encoding over these batches, we can apply gradient accumulation to trade runtime for reduced memory consumption. We thus analyze PERK across varying gradient accumulation steps, ranging from 2 to 16. Note that for the minimum context length of 1024 tokens, which yields a batch of 8 sequences, gradient accumulation with 16 steps is infeasible as it exceeds the batch size.
PERK scales more efficiently in memory and runtime on extremely long contexts.
Figure 6a shows that, as the training context length grows from 1K to 8K tokens, Reckoning quickly runs into Out-Of-Memory (OOM) errors at 2K, while PERK with increasing truncation steps continues to fit the maximum GPU memory. We successfully validate PERK’s stronger scalability on long contexts.
Figure 6b for inference complexity shows that PERK provides more efficient memory and runtime scaling for extremely long contexts compared to FT-ICR. While FT-ICR is initially more efficient, its memory and runtime grow rapidly, leading to OOM errors at a context length of 128K. In contrast, PERK can manage the long sequences through gradient accumulation, which, while increasing runtime, reduces the memory footprint. For instance, at 8K tokens, increasing the accumulation steps from 1 to 16 reduces memory usage from 35.2GB to 5.9GB, although the runtime increases from 1.9s to 8.5s. Crucially, even with this runtime trade-off, PERK runs faster than FT-ICR at very long contexts where both can operate. At 64K tokens, FT-ICR takes 32.6s and 55.7GB, whereas PERK with 16 steps of accumulation completes in 11.4s, using only 19.6GB. With 8 steps, it is even faster at 9.7s, using 35.2 GB. Ultimately, at 128K tokens, where FT-ICR fails, PERK with 16 steps successfully processes the context using 35.2GB in 20.9s, showing that PERK provides a practical path to handle extreme context lengths efficiently in both memory and runtime when compared to standard approaches.
6 Related Work
Test-Time Learning
MAML [17] introduces optimization-based model-agnostic meta-learning for few-shot adaptation. [13] applies MAML to GPT models, integrating knowledge into model parameters through test-time gradient updates. Some methods achieve test-time learning in the model architecture. [58] proposes RNN layers with learnable hidden states updated via self-supervised gradients on test inputs. Titans [8] incorporates differentiable memory with transformers for rapid inference-time memorization. Building on the Titan memory, ATLAS [7] proposes a new transformer architecture that contains the differentiable memory module as part of the network to support large-scale distributed pretraining. Other approaches adapt through context alone. MetaICL [43] meta-learns from in-context examples without gradient updates. Our method, PERK, introduces parameter-efficient meta-learning using lightweight adapters, improving scalability for long contexts.
Long-Context Language Models
Developing long-context models has become a major research topic of its own. Many works explore extending transformer-based language models’ context windows with minimal training, either by position extrapolation [12, 47, 63, 69] or improving the transformer’s attention mechanism [9, 44, 56, 27]. Other works show that using the default attention, applying simple position extrapolation, and fine-tuning the model on long documents [18, 42] or synthetically-generated long data [1, 64] yields much stronger results. The FT-ICR baseline in our work follows this paradigm. Other LLM works [37, 20, 19] propose achieving long-context capabilities through a long-context continued training stage between standard pre-training and supervised fine-tuning. In contrast to all the approaches above, PERK achieves long-context reasoning through the ability to encode long sequences using test-time gradient updates on a lightweight model adapter.
Long-Context Architectures
In parallel, there have been many efforts in designing more efficient architectures, for example, linear attention with RNNs [11, 46, 21, 15, 66, 6] and alternative attention architectures [53, 59, 22]. However, these new architectures often require training from scratch, and many have inherent limitations in terms of long-context recall [26, 3]. Recent works also explore hybrid models [37, 16] or distilling existing LLMs into hybrid models [61] and show promising results. In comparison, PERK directly augments off-the-shelf pre-trained LLMs without requiring additional architecture modifications or extensive pre-training, while naturally being parameter-efficient.
Long-Context Evaluation
Many benchmarks have been proposed to evaluate language models’ long-context abilities [31, 55, 4, 71, 23, 70, 68, 62]. Some works focus on the impact of extending input lengths with irrelevant contexts [32, 35] or changing relevant information positions [39] on LLM reasoning. Many works also reframe natural-language-processing tasks, such as retrieval [34, 49], summarization [30], and in-context learning [36, 65, 10], as long-context challenges. In this work, we define Drops-in-the-Ocean as a new long-context evaluation paradigm that makes the long context as distributionally similar to relevant information as possible.
7 Conclusion
In this work, we introduce PERK, a parameter-efficient meta-learning approach that learns to perform long-context reasoning by encoding information into a lightweight adapter through test-time gradient updates. Our experiments on multiple long-context benchmarks (Needle-in-the-Haystack, Drops-in-the-Ocean, and API retrieval) demonstrate substantial performance gains over models finetuned for classical in-context reasoning across reasoning complexity and context lengths. Our generalization analysis highlights PERK’s strong test-time length extrapolation ability, where the model can generalize to unseen contexts 32 times longer than the training data. When relevant information shifts positions at test-time, PERK robustly generalizes across different positions. Finally, we show that PERK scales efficiently at inference to extremely long contexts. Overall, PERK excels in performance, generalization, and efficiency for long-context reasoning.
Appendix A Datasets
In this section, we provide details on the datasets used for each task in our experiments and analysis. For each dataset, we visualize an example in Figure 7.
A.1 BabiLong
We construct the train and evaluation data using the data generation framework proposed by [32]. As we mentioned in the experiment setup, we mainly focus on tasks involving single-hop (QA1), two-hop (QA2), and three-hop (QA3) reasoning.
Tasks
The main task for BabiLong is to reason over facts distributed in a long context. For each QA task, we have a set of facts, including distractors and supporting facts. Each fact describes an agent’s action or movement (which also describes the agent’s location). For example, Mary journeyed to the garden is a movement, where Mary is the agent, and garden is Mary’s current location. As an example for actions, John picked up the apple there describes John’s action of picking up an apple. Together with multiple agents performing multiple movements and actions, the set of facts became a temporal sequence of events.
QA1 focuses on one-hop reasoning, where the problems can be answered by using a single supporting fact. For example, in a sequence of actions and movements, the agent John’s last movement is John moved to the bedroom. Then the one-hop question is “Where is John located now?”. By the last movement alone, we can find the answer “bedroom”.
QA2 focuses on one-hop reasoning, where the problems can be answered by using two supporting facts, where one fact is a movement and the other one is an action. For example, in the event sequence, we have John moved to the garden and John dropped the apple there. Using these two facts, we can answer “Where is the apple now?”, which is “garden”.
QA3 focuses on one-hop reasoning, where the problems can be answered by using three supporting facts, as a combination of movements and actions. For example, we have three supporting facts: Daniel went back to the office, Daniel went back to the bedroom, and Daniel left the milk in order. We can infer the answer to the question “Where was the milk before the bedroom?” is “office”.
Build Long Context
To build the long context, we use the framework to construct the tasks with long natural documents sampled from FineWeb-edu [41], specifically from the CC-MAIN-2024-42 subset, ensuring the content postdates our base model’s knowledge cutoff. Given a particular context window, such as 8192, we sample that many tokens from the corpus as the Haystack long context.
We need to convert a long context into a batch of subsequences, such that PERK can encode them in parallel through test-time gradient-based adaptation. In practice, we set the effective context length (i.e., the number of tokens of a subsequence on average) as 256. We divide the long context into chunks where each chunk is approximately 256 tokens long. Then, we randomly distribute the set of facts into these chunks. For each fact, we randomly select a chunk and then insert this fact into the chunk at a random position. To preserve the temporal order of the facts (since they come from a sequence of events with temporal order), we index each fact with a prefix. Finally, we have a batch of subsequences where each sequence might contain some facts at random positions.
A.2 Student Records
In the main paper, we propose Drops-in-the-Ocean (DIO), a new evaluation setting where long contexts are formed from structurally similar documents. We aim to create problems in which the relevant and irrelevant information will likely be distributionally similar, making it more difficult to identify critical facts. We construct a synthetic dataset, Student Records, where each context simulates a database containing multiple student records. Each record includes several attributes: ID, name, school, major, and grade.
Tasks
The task operates on top of a database that consists of multiple records. Each record holds information about a student. The record is structured as follows:
Student Id: 3109998, Student Name: Alison Keith, Year: Sophomore
School: School of Computer Information and Data Sciences, Major: Computer Science, Grade: 72
We generate the student IDs and grades using a random number generator. The student names are generated from the Name package of Python. We define a pool of years, schools, and courses. Then, we randomly select each entity for each student. Crucially, to test generalization, specific entities (e.g., IDs, names) and attribute value combinations encountered in the test set are disjoint from those used during training.
Based on the database of student records, we construct three main tasks: Recall (retrieving attributes for a specific ID), Relation (comparing attributes between two IDs), and Aggregate (calculating the maximum, minimum, and average grade across all students).
Recall tests a model’s ability to retrieve an attribute of a student given a specific student ID number. For example, we can ask “What major does student 1778924 study?”. The model should answer “What major does student 1778924 study?”. The questions cover all attributes for testing coverage.
Relation tests a model’s ability to compare the attributes of two students. For example, for the attribute of grade, we can ask “Does student 1778924 have a higher grade than student 7077015?”. The model should be able to compare their grades and infer “Yes”. Or we can also ask “Do student 1778924 and student 7077015 study the same major?”, the model should recognize that their majors are different and answer “No”.
Aggregation is the most difficult task in this dataset. This task requires many-hop reasoning ability, where the model needs to aggregate attributes from all students in a long context to make an accurate conclusion. We focus on three types of numerical aggregation among the students’ grades: (1) Maximization, (2) Minimization, and (3) Averaging. For example, we ask the model “What is the average grade of all students”, the model should be able to compute the sum of all student grades, count the number of total students in the context, and then compute the average, “49.3”.
Figure 7: Dataset Examples. Here we show an example from the three datasets we used: (Top) BabiLong, (Middle) Student Records, and (Bottom) Lost-in-the-Middle API Retrieval.
Build Long Context
Context length scales directly with the number of student records included. For a given context window, we continuously append student records to the context until the number of tokens exceeds the context window. To convert the long context into a batch of subsequences for PERK, we simply treat each student record as an individual subsequence in the batch, since they are already independent documents by design.
A.3 Lost-in-the-Middle API Retrieval
To analyze a model’s robustness against relevant information positions, we follow the setting proposed by Lost in the Middle [39]. We use API documents from [45], where each document has an associated API call, and create long contexts containing multiple API-call-to-document pairs. We place the relevant API document in a particular user instruction at different locations in the context (beginning, middle, end, and random position).
Task
We formulate the main task as gold API retrieval. In this task, the model must retrieve the correct API call based on user instructions and documents. Using fixed 4k and 8k token contexts, we train PERK and FT-ICR models on contexts with the target document at varying positions and test them on contexts distributed across all positions.
Build Long Context
We follow Student Records and scale the context length with the number of API documents included. For a given context window, we continuously append API documents to the context until the number of tokens exceeds the context window. Similarly, since each API document is independent, we treat a document as a subsequence in the batch for PERK. Since the documents are processed in parallel as a batch, they do not preserve any ordering information. Thus, by design, PERK is robust to the position of the relevant information in a context.
Appendix B Training and Inference Details
Hardware Information
All training runs are conducted on a single Nvidia H100 GPU with 93GB of VRAM. The training jobs run on an internal research cluster managed by Slurm. Each job is provided with 64 CPUs and 64GB of RAM. The GPU is completely utilized without sharing with other concurrent jobs.
Standard Hyperparameters
Each training (for PERK and baselines) sets 2 epochs as the maximum runtime with early stopping based on validation loss. In practice, most training runs typically stop early, within one epoch of runtime. We set the baseline training and PERK’s outer loop learning rate to 1e-5, and the weight decay to 0.01. We use a cosine learning rate scheduler with 0.03 of the total optimization steps as the warm-up phase. We use PyTorch Lightning  [^1] https://lightning.ai/docs/pytorch/stable/ as the trainer for PERK. The Qwen and GPT-2 FT-ICR baselines are trained using the Open-Instruct [33] trainer from AI2, and the Mamba FT-ICR baselines are trained using the Huggingface Trainer  [^2] https://huggingface.co/docs/transformers/en/main_classes/trainer. For the stress test of length generalization beyond 64K tokens, we change the default maximum position of 32768 to 131072.
LoRA Hyperparameters
We set LoRA rank to 256 for both PERK GPT-2 and PERK Qwen. We use the default values for the other hyperparameters, including setting the alpha to 16 and the dropout to 0.1. We apply LoRA to all modules in the model architecture. For GPT-2, the number of active training parameters of PERK is 38M, 30% of the total parameter number. For Qwen, the number of active training parameters of PERK is 140M, which is also 30% of the total parameter number. We set the scaling factor $\alpha$ to 16 initially and enable rank-stabilized LoRA fine-tuning [28], which results in a target $\alpha$ of 256, corresponding to a rank of 256.
Meta-Learning Hyperparameters
For the inner loop adaptation, we use 4 inner loop gradient steps. For the truncated gradient unrolling, we truncate steps 1 and 2 for input context length $\leq$4K, and we truncate steps 1, 2, and 3 for context length $>$ 8K. We follow [67]’s practice on how to train a good MAML model and use per-layer-per-step dynamic learning rate for the inner loop, using 5e-5 as the initial value. For the language modeling loss in the inner loop, we use a weighted version similar to CaMeLS [25]. The token-level weights are predicted via a weighting model implemented as a two-layer MLP network. The parameters of the weighting model are co-meta-learned along with the LoRA parameters and the adaptive per-layer-per-step learning rates. We use the differentiable version of the AdamW [40] optimizer for the inner loop optimization, implemented by the Higher  [^3] https://github.com/facebookresearch/higher library for higher-order optimization. We compute higher-order derivatives using the automatic differentiation mechanism implemented by PyTorch [^4] https://docs.pytorch.org/tutorials/beginner/blitz/autograd_tutorial.html. The outer loop optimizer is the standard PyTorch implementation of AdamW.
Evaluation Hyperparameters
At inference time, we generate the answer for a reasoning problem via greedy decoding (temperature is set to 0). The maximum number of tokens that can be generated is 512. The EOS token is set to the default one defined in the tokenizer. The generation is handled by the text-generation pipeline implemented by Huggingface  [^5] https://huggingface.co/docs/transformers/v4.43.4/en/main_classes/pipelines.
Appendix C Additional Analysis
Method QA1 QA2 QA3
FT-ICR 91.1 70.5 48.6
PERK LoRA-16 99.0 96.9 84.7
PERK LoRA-256 100 98.1 89.1
Table 1: Analyzing LoRA rank’s impact on PERK’s performance.
C.1 Impact of LoRA Rank in PERK
In this section, we analyze the impact of LoRA’s rank on PERK’s long-context reasoning performance. We focus on the BabiLong tasks with a 1024 context window for an efficient ablation study. We apply PERK to Qwen-2.5-0.5B with LoRA rank of 16 and 256 and evaluate on the BabiLong tasks. We show the results in Table 1. We also include the performance from the FT-ICR Qwen baseline as a reference. The results show that PERK’s performance only decreases slightly when the rank is reduced to 16. However, PERK still maintains a much stronger performance than the FT-ICR Qwen baseline. The analysis shows that PERK can maintain its strong performance with a more parameter-efficient setting, highlighting its potential for more efficient long-context reasoning.
Method QA1 QA2 QA3
FT-ICR 91.1 70.5 48.6
PERK 2 Steps 99.2 94.5 82.4
PERK 4 Steps 100 98.1 89.1
Table 2: Analyzing the impact of the number of inner loop adaptation steps on PERK’s performance.
C.2 Inner Loop Adaptation Steps
In this section, we analyze the impact of inner loop adaptation steps on PERK’s long-context reasoning performance. Similarly, we focus on the BabiLong tasks with a 1024 context window. We again apply PERK to Qwen-2.5-0.5B for 2 and 4 inner loop steps. With 4 steps, we truncate the first 2 steps in backpropagation. With 2 steps, we truncate the first step in backpropagation. We show the results in Table 2. The results show that reducing the number of inner loop adaptation steps decreases the performance more noticeably than reducing LoRA rank. However, PERK still outperforms FT-ICR Qwen on all three tasks. The results here suggest that we should select a higher number of inner loop adaptation steps to maximize the performance. If in settings where GPU memory might be limited, reducing the number of inner loop steps can reduce the performance, but it is still able to outperform the baseline. However, the analysis is done with limited hardware resources, which limits our analysis to 4 steps. To further analyze the impact of the inner loop step counts, more systematic experiments in a scalable hardware setting are needed for future work.